# Tracer diffusion beyond Gaussian behavior: explicit results for general single-file systems

### v1

@@ Abstract
MFT has proved to be useful to study a wide range of observables, including even a microscopic observable such as the position $X_{T}$ of a single tracer, for which different approaches have been devised. (i) The first one relies on expressing $X_{T}$ as a functional of the density of particles $X_{T}=X_{T}[\rho]$ [18, 19]. This can be done since the tracer effectively “cuts” the system into two part, in which the number of particles is conserved due to the noncrossing condition. In practice, this is however tricky due to the emergence of discontinuities in the density profiles at the position of the tracer [18, 19]. (ii) A second approach that circumvents this issue consists in introducing a generalised current [21, 22] defined as the number of particles crossing a fictitious moving wall. The tracer is then located at the position where this current vanishes, again due to the noncrossing condition. (iii) An alternative method, which we apply here, consists in using a mapping between different single-file systems, in which the position $X_{T}$ of the tracer in the original model is mapped onto (the opposite of) the integrated current $\tilde{Q}_{T}$ through the origin in a dual model. More precisely, the current is defined from the density $\tilde{\rho}(x,t)$ in the dual model as
@@ I.2 Mapping onto a dual problem of current
The idea behind this transformation is that a given single-file system can be viewed in two equivalent ways: (i) either by looking at the positions of particles, and thus their density $\rho(x,t)$ (ii) or by looking at the distances between the particles. In this second point of view, the variable $y$ represents the ”label” of the particle located at position $x$ at time $t$, while $\tilde{\rho}(y,t)$ is the distance between this particle and the next one. Although equivalent, these two points of view give rise to two different models of single-file systems, with different transport coefficients related by (S6).
@@ II.2 Second order
$Q_{2}(y,t)=-\frac{e^{-\frac{2y^{2}}{t}}}{2\pi}-\frac{e^{-\frac{y^{2}}{t}}}{8\pi\sqrt{t}}+\frac{e^{-\frac{y^{2}}{t}-\frac{y^{2}}{1-t}}}{8\pi}\sqrt{\frac{1-t}{t}}\\ +\frac{y\>e^{-\frac{y^{2}}{t}}}{8\sqrt{\pi\>t}}\left[\operatorname{erfc}\left(\frac{y}{\sqrt{t}}\right)-\operatorname{erfc}\left(\frac{y}{\sqrt{1-t}}\right)\right]+\frac{1}{4}\mathrm{T}\left(y\sqrt{\frac{2}{t}},\sqrt{\frac{t}{1-t}}\right)\>.$ (S30)
@@ III.2 Correlation profiles
where $f$ is the function (S48). Finally, Eqs. (S67-S69) give the expression of the bath-tracer correlation profiles up to order $3$.
@@ III.5 A conjecture for the cumulant generating function
The two equations (S79,S81) are the boundary conditions given in the main text. They have first been derived for the specific case of the SEP, and it has recently been shown that they hold for any single-file model with arbitrary $D(\rho)$ and $\sigma(\rho)$ [13].
Although the two boundary conditions (S79,S81) are very general, and take a compact physical form, they give less information than the three equations (S78). In particular, Eqs. (S79,S81) do not involve the cumulant generating function $\hat{\psi}$, while (S78) does. To fully generalise (S78) to any $D(\rho)$ and $\sigma(\rho)$, we need the last relation that relates $\hat{\psi}$ and $\Phi(0^{\pm})$, $\Phi^{\prime}(0^{\pm})$. This relation was very useful in [11, 5, 6] as it allows to easily compute $\hat{\psi}$ from $\Phi$. To obtain this relation for any single-file system, let us start by rewriting (S78) as
@@
Finally, from our explicit results on $\kappa_{4}$ (S62) and the profiles (S67-S69), we check that (S84) holds, at least up to order $4$ in $\lambda$ included.

### v2

@@ Abstract
MFT has proved to be useful to study a wide range of observables, including even a microscopic observable such as the position $X_{T}$ of a single tracer, for which different approaches have been devised. (i) The first one relies on expressing $X_{T}$ as a functional of the density of particles $X_{T}=X_{T}[\rho]$ [18, 19]. This can be done since the tracer effectively “cuts” the system into two part, in which the number of particles is conserved due to the noncrossing condition. In practice, this is however tricky due to the emergence of discontinuities in the density profiles at the position of the tracer [18, 19]. (ii) A second approach that circumvents this issue consists in introducing a generalised current [21, 22] defined as the number of particles crossing a fictitious moving wall. The tracer is then located at the position where this current vanishes, again due to the noncrossing condition. (iii) An alternative method, which we apply here, consists in using a mapping between different single-file systems, in which the position $X_{T}$ of the tracer in the original model is mapped onto (the opposite of) the integrated current $\tilde{Q}_{T}$ through the origin in a dual model (see Fig. 1). More precisely, the current is defined from the density $\tilde{\rho}(x,t)$ in the dual model as
@@ I.2 Mapping onto a dual problem of current
This is a generalisation of previously known mappings between specific models of single-file systems, such as between the SEP and the zero range process [4, 5], and between the random average process (RAP) and the Kipnis Marchioro Presutti (KMP) model [6]. The idea behind this transformation is that a given single-file system can be viewed in two equivalent ways: (i) either by looking at the positions of particles, and thus their density $\rho(x,t)$ (ii) or by looking at the distances between the particles. In this second point of view, the variable $y$ represents the ”label” of the particle located at position $x$ at time $t$, while $\tilde{\rho}(y,t)$ is the distance between this particle and the next one. Although equivalent, these two points of view give rise to two different models of single-file systems, with different transport coefficients related by (S6).
@@ II.2 Second order
$Q_{2}(y,t)=-\frac{e^{-\frac{2y^{2}}{t}}}{2\pi}-\frac{e^{-\frac{y^{2}}{t}}}{8\pi\sqrt{t}}+\frac{e^{-\frac{y^{2}}{t}-\frac{y^{2}}{1-t}}}{8\pi}\sqrt{\frac{1-t}{t}}\\ +\frac{y\>e^{-\frac{y^{2}}{t}}}{8\sqrt{\pi\>t}}\left[\operatorname{erfc}\left(\frac{y}{\sqrt{t}}\right)-\operatorname{erfc}\left(\frac{y}{\sqrt{1-t}}\right)\right]+\frac{1}{4}\mathrm{T}\left(y\sqrt{\frac{2}{t}},\sqrt{\frac{t}{1-t}}\right)\>,$ (S30)
where $T$ is the Owen-T function, defined by [8]
$T(h,a)=\frac{1}{2\pi}\int_{0}^{a}\frac{e^{-\frac{h^{2}}{2}(1+x^{2})}}{1+x^{2}}\mathrm{d}x\>.$ (S31)
@@ III.1.2 Recovering known results in specific cases
We briefly describe the models listed in this table. In the simple exclusion process (SEP), each particle can hop onto a neighbouring site if it is empty. The random average process describes particles on a continuous line which can hop to a random fraction of the distance to a neighbouring particle. The Kipnis-Marchioro-Presutti describes a lattice in which each site hosts a continuous variable (a mass). At random times the total mass of two neighbouring sites is randomly redistributed between them. The hard Brownian particles is a model of hardcore particles performing Brownian motion, with the constraint that they remain in the same order. The hard rod gas is similar, but the particles have a finite length $\ell$ and cannot overlap. Finally, the double exclusion process is similar to the SEP, but the particles occupy two neighbouring sites and can only hop by one unit if the site in the randomly chosen direction is empty.
@@ III.2 Correlation profiles
where $f$ is the function (S49). Finally, Eqs. (S68-S70) give the expression of the bath-tracer correlation profiles up to order $3$.
@@ III.5 A conjecture for the cumulant generating function
The two equations (S80,S82) are the boundary conditions given in the main text. They have first been derived for the specific case of the SEP, and it has recently been shown that they hold for any single-file model with arbitrary $D(\rho)$ and $\sigma(\rho)$ [15].
Although the two boundary conditions (S80,S82) are very general, and take a compact physical form, they give less information than the three equations (S79). In particular, Eqs. (S80,S82) do not involve the cumulant generating function $\hat{\psi}$, while (S79) does. To fully generalise (S79) to any $D(\rho)$ and $\sigma(\rho)$, we need the last relation that relates $\hat{\psi}$ and $\Phi(0^{\pm})$, $\Phi^{\prime}(0^{\pm})$. This relation was very useful in [13, 9, 10] as it allows to easily compute $\hat{\psi}$ from $\Phi$. To obtain this relation for any single-file system, let us start by rewriting (S79) as
@@
Finally, from our explicit results on $\kappa_{4}$ (S63) and the profiles (S68-S70), we check that (S85) holds, at least up to order $4$ in $\lambda$ included.